A formula for calculating the magnitude of an earthquake is M=2/3log(EE0)

 that uses the common (base 10) logarithm. This is called the Moment Magnitude Scale (MMS), an alternative to the more well known Richter Scale. One earthquake has magnitude 3.9

 on the MMS. If a second earthquake has 800

 times as much energy as the first, find the magnitude of the second quake.

 



1
Expert's answer
2021-01-26T04:34:00-0500

"M= \frac{2}{3log(EE')}"


when M= 3.9,


3.9= "\frac {2}{3log(EE')}" [ log (base 10) is used ]


log(EE')= "\frac{2}{3*3.9}"


log(EE')= 0.171


EE'= 100.171 =1.482


when energy became 800 times as much as energy as first earthquake had, then


"M= \frac{2}{3log(800*1.482)}"


"M= \frac{2}{3*3.074}\\M= 0.217"


hence, the magnitude of second earthquake is M= 0.217